x+1/x=12,then the value of x-1/x=? please reply me.
Answer by Alan3354(69443)

x+1/x=12,then the value of x-1/x=? please reply me.
---------------
x+1/x=12
Is it (x+1)/x ? or
x + (1/x) ?
Parentheses are free, use some.